# CCTP-2025-7 QMUL-PH-25-13 ITCP-IPP-2025/7

### v1

@@ 1 Introduction
Foundation models [1], like the flagship Large Language Models (LLMs) GPT-4 of Open-AI [2], and the Llama models [3, 4, 5] of Meta, exhibit remarkable performance on a wide range of tasks (from logical reasoning to creative writing), gaining a lot of traction in recent years over diverse domains, with both commercial and scientific applications. [^1] For a more complete list of popular LLMs, the reader can consult, for example, the blog post https://explodingtopics.com/blog/list-of-llms and the GitHub page https://github.com/Hannibal046/Awesome-LLM. For the scientific community, in particular, the ability of LLMs to accelerate the scientific process and discovery is a topic of obvious interest. For example, a recent comprehensive survey of over 260 scientific LLMs in Mathematics and the Natural Sciences [6] summarizes applications that range from the generation of novel scientific ideas to problem-solving in Mathematics, Physics, Chemistry and Materials Science, Computer Science, Biology and Medicine, and Environmental Sciences.
@@
Second, we believe that Theoretical Physics presents rich opportunities for future developments in AI. On the one hand, it lies close to pure Mathematics: it is heavily based on mathematical language and reasoning; it values mathematical rigor and consistency. On the other hand, Theoretical Physics, being based on natural phenomena, also involves physics-based hypotheses, reasoning and consistency checks which do not directly follow from mathematics alone (e.g. one cannot arrive at the principle of wave-particle duality or Schrodinger’s wave-equation via mathematics alone). While a significant push is being made to equip LLMs with mathematical reasoning capabilities, we believe a similar effort in teaching physics-based reasoning and logic to LLMs will be help them build a better internal model of the world.
@@ 2 Datasets and Training
The fine-tuning process was carried out using the Hugging Face ecosystem. In order to improve resource efficiency, the Llama model weights were quantized to 4-bit precision with the bitsandbytes package [16, 17] and we employed LoRA [18] to reduce the number of trainable parameters. As mentioned in the Introduction, we experimented with two versions of LoRA: one applied to the QKV matrices only (LoRA-QKV), and one applied to all projection matrices (LoRA-all). Our complete hyperparameter list is given in Table 2.
@@
At this point we would like to pause and discuss a potential ambiguity in the definition of perplexity. While the definition of perplexity for one sequence is the exponential of the cross entropy loss, it is not a priori clear how the perplexity of a dataset, consisting of multiple sequences, is defined. We have seen instances that seem to suggest that one should take the geometric mean [20]. However, we are of the opinion that the more relevant metric is the arithmetic mean; see Appendix B for a related discussion. In this paper when we discuss perplexity over a dataset, we mean the arithmetic mean of the perplexities of the individual sequences.
@@ 3 Evaluation Metrics
To evaluate the performance of the fine-tuned models, we processed each abstract in the target hep-th test dataset (denoted s1) by splitting it into a prompt and ground truth as follows. Each abstract was first divided into sentences, with the total number of sentences in an abstract denoted by $N$ [^5] If an abstract contained only a single sentence, it was instead split into $N$ words. The prompt was constructed by taking the first $\lceil\frac{N}{2}\rceil$ sentences, while the remaining sentences formed the ground truth. The prompts were then run through the models fine-tuned on the datasets of Table 1—which we will refer from here on by the same abbreviation s1-s10—to obtain completions with temperature 0.8 and a maximum of 1024 tokens. We point out that during training we quantized the base LLaMA model weights to 4-bit precision and fine-tuned 16-bit LoRA adapters on top. However, for evaluation we merged the trained LoRA weights back into the base model and loaded the entire merged model in 16-bit precision.
@@ 3.1 Perplexity
For each of our fine-tuned models, including the LoRA-QKV and LoRA-all variants, we evaluated the mean batch perplexity on the hep-th (s1) test dataset. In order to get a comparative sense between models, we also computed the mean and standard deviations of these perplexities using the bootstrap technique. In this technique the sampling distribution of a statistic is estimated by resampling with replacement from the original dataset. Statistics are then extracted by computing the arithmetic mean and standard deviation of the bootstrap distribution. The results are displayed in Table 3, with a visual presentation given in Figure 3.3.
Unsurprisingly, all LoRA-all models fine-tuned on datasets including hep-th abstracts (s1, s3-s8, s10) exhibited lower perplexity compared to the base model. On the contrary, models that did not include hep-th abstracts (s2, s9) had similar or higher perplexity. We also noted that the overall size of the dataset matters, with the lowest perplexity variants being those that included all hep-th samples augmented with other datasets (hep-ph, gr-qc, cs, q-bio).
@@
Llama 11.20 11.20
s1 10.32 10.04
s2 11.38 10.62
s3 9.83 10.02
s4 10.41 10.08
s5 10.35 10.07
s6 10.34 10.05
s7 10.00 10.04
s8 9.90 10.04
s9 12.16 10.64
s10 9.93 10.00
Table 3: Comparison of perplexity means across base and fine-tuned model configurations. The perplexity of the base Llama model without any fine-tuning (LoRA-all or LoRA-QKV) has been included to make the comparison to the baseline easier.
@@ 3.2 Semantic Similarity
On average we observed that the base model scored $0.88\pm 0.08$ while all the fine-tuned models (LoRA-QKV and LoRa-all) $0.90\pm 0.07$. Therefore, from the perspective of SemScore all models (base and fine-tuned) exhibit comparable performance in semantic similarity compared to the ground truth (with the fine-tuned models perhaps having a marginal edge). We note that the minimum values of the cosine similarities exhibited more pronounced variation: for the base model the minimum was 0.007, for the LoRA-all models the minimum was above 0.1 (with the exception of the s2 model, that was not trained on hep-th, at 0.06), and for the LoRA-QKV models the minimum had a much higher variation ranging from $-0.015$ to 0.15. For example, the s10 LoRA-QKV model (trained on hep-th, q-bio, cs) had minimum cosine similarity at $-0.001$.
@@ 3.3 Human evaluation
In an effort to evaluate the quality of the completions of the fine-tuned models from an end-user perspective, we independently went through tens of completions for the base and fine-tuned models, examining and rating them on the basis of the following criteria: grammatical and stylistic form, logical coherence, relevance, scientific accuracy and creativity/originality. Each of us inspected independently the completions on 5 of their own past scientific papers and 5 randomly chosen, famous papers in the hep-th category. In addition, we asked the (free version of the) popular commercial LLMs ChatGPT, Claude, Gemini and DeepSeek to complete the same abstracts using the same prompt. This allowed us to make a first qualitative comparison with full-fledged LLMs of vastly different size that have been trained on much larger datasets.
Some notable observations are as follows:
@@
The fine-tuned models could use technical language and, although training exclusively on abstracts limited their factual accuracy, they demonstrated competitive performance relative to commercial LLMs. The example in Tables 4 demonstrates how two of our fine-tuned models (s3 and s8 LoRA-all) performed against the ground truth and ChatGPT-o. The fine-tuned models can clearly use technical language in sensible and relevant contexts, but are not 100% factually correct. For instance, in this example, the s3 model sensibly refers to the class-$\mathcal{S}$ $N=2$ superconformal field theories, which arise in compactifications of M5-branes, while the s8 model similarly picks up on the related AGT correspondence. The completion of ChatGPT, on the other hand, is very close to the content of the actual paper referencing key components of the proof. The completion of the base Llama model essentially repeats elements of the prompt, as well as a line of (false) metadata at the end.
@@
In App. C we further discuss some of these observations with a collection of a few additional explicit completions.
@@ Appendix A Data Curation
[…]
@@
Appendix B Cross Entropy Loss and Perplexity
@@
$\mathcal{L}=-\frac{1}{T}\sum_{t=0}^{T}\sum_{i\in|\mathcal{V}|}y_{t+1,i}\log(p_{\theta}(w_{t+1,i}|w_{1:t}))\;.$ (B.2)
@@
$\mathcal{P}=\Bigg{(}\prod_{t=0}^{T}\frac{1}{p_{\theta}(w_{t+1,\text{Id}(y_{t+1})}|w_{1:t})}\Bigg{)}^{1/T}\;,$ (B.3)
@@
The fact that all three ended up with the same perplexity of 3 is purely stochastic. Let us see what happens when we do sentence completion with the context length set to two words:
@@
What we just discussed is the ground truth perplexity coming from the fact that the training data itself has ambiguity in how a sequence is completed having started off with "I". This is the lowest perplexity achievable. Models in general will have higher perplexity. The normalization is done so as to be able to compare perplexities across sequences of different lengths.
The above discussion involved the perplexity evaluated against one ground truth sequence. How do we compute it for a set of sequences? A naive answer, commonly used in the literature as it is computationally simpler, would be to take the cross entropy loss for the whole dataset and exponentiate it. This would amount to a geometric mean of perplexities of sequences. Another possibility is to take the arithmetic mean of perplexities of the sequences as this would give the average number of branching factors of the language model given a prompt.
@@
Appendix C Additional examples of completions

### v2

# arXiv:2508.03716v2[cs.CL] 27 Feb 2026

@@ 1 Introduction
Foundation models [1], like the flagship Large Language Models (LLMs) GPT-4 of Open-AI [2], and the Llama models [3, 4, 5] of Meta, exhibit remarkable performance on a wide range of tasks (from logical reasoning to creative writing), gaining a lot of traction in recent years over diverse domains, with both commercial and scientific applications. [^1] For a more complete list of popular LLMs, the reader can consult, for example, the blog post https://explodingtopics.com/blog/list-of-llms and the GitHub page https://github.com/Hannibal046/Awesome-LLM. For the scientific community, in particular, the ability of LLMs to accelerate the scientific process and discovery is a topic of obvious interest. For example, a comprehensive survey of over 260 scientific LLMs in Mathematics and the Natural Sciences [6] summarizes applications that range from the generation of novel scientific ideas to problem-solving in Mathematics, Physics, Chemistry and Materials Science, Computer Science, Biology and Medicine, and Environmental Sciences; see also [7] for very recent results in this direction, where some new mathematical results were obtained with the use of ChatGPT-5.
@@
Second, we believe that Theoretical Physics presents rich opportunities for future developments in AI. On the one hand, it lies close to pure Mathematics: it is heavily based on mathematical language and reasoning; it values mathematical rigor and consistency. On the other hand, Theoretical Physics, being based on natural phenomena, also involves physics-based hypotheses, reasoning and consistency checks which do not directly follow from mathematics alone (e.g. one cannot arrive at the principle of wave-particle duality or Schrodinger’s wave-equation via mathematics alone). While a significant push is being made to equip LLMs with mathematical reasoning capabilities, we believe a similar effort in teaching physics-based reasoning and logic to LLMs will help them build a better internal model of the world.
@@ 2 Datasets and Training
The fine-tuning process was carried out using the Hugging Face ecosystem. In order to improve resource efficiency, the Llama model weights were quantized to 4-bit precision with the bitsandbytes package [17, 18] and we employed LoRA [19] to reduce the number of trainable parameters. LoRA works by freezing the pre-trained model weights and inserting trainable low-rank matrices into each layer, allowing efficient adaptation with a small fraction of the original parameters—–this is analogous to treating the fine-tuning process as a perturbation on the base model’s weight matrices. As mentioned in the Introduction, we experimented with two versions of LoRA: one applied to the QKV matrices only (LoRA-QKV), and one applied to all projection matrices (LoRA-all). A detailed hyperparameter list is given in Table 2, with the complete settings presented in Appendix B.
@@
At this point we would like to pause and discuss a potential ambiguity in the definition of perplexity. While the definition of perplexity for one sequence is the exponential of the cross entropy loss, it is not a priori clear how the perplexity of a dataset, consisting of multiple sequences, is defined. We have seen instances that seem to suggest that one should take the geometric mean [21]. However, we are of the opinion that the more relevant metric is the arithmetic mean; see Appendix C for a related discussion. In this paper when we discuss perplexity over a dataset, we mean the arithmetic mean of the perplexities of the individual sequences.
@@ 3 Evaluation Metrics
To evaluate the performance of the fine-tuned models, we processed each abstract in the target hep-th test dataset (denoted s1) by splitting it into a prompt and ground truth as follows. Each abstract was first divided into sentences, with the total number of sentences in an abstract denoted by NN. [^5] If an abstract contained only a single sentence, it was instead split into $N$ words. The prompt was constructed by taking the first $\lceil\frac{N}{2}\rceil$ sentences, while the remaining sentences formed the ground truth. The prompts were then run through the models fine-tuned on the datasets of Table 1—which we will refer from here on by the same abbreviation s1-s10—to obtain completions with temperature 0.8 and a maximum of 1024 tokens. We point out that during training we quantized the base LLaMA model weights to 4-bit precision and fine-tuned 16-bit LoRA adapters on top. However, for evaluation we merged the trained LoRA weights back into the base model and loaded the entire merged model in 16-bit precision. The full settings used for inference can be found in Appendix B.
@@ 3.1 Perplexity
For each of our fine-tuned models, including the LoRA-QKV and LoRA-all variants, we evaluated the mean batch perplexity on the hep-th (s1) test dataset. In order to get a comparative sense between models, we also computed the mean and standard deviations of these perplexities using the bootstrap technique. In this technique the sampling distribution of a statistic is estimated by resampling with replacement from the original dataset. Statistics are then extracted by computing the arithmetic mean and standard deviation of the bootstrap distribution. The results are displayed in Table 3, with a visual presentation given in Figure 3.3. For reference, Transformer-XL achieved 18.9 perplexity on WikiText-103 [22]. Although our fine-tuned models performed better on hep-th abstracts, direct comparison is complicated by differences in tokenizer vocabulary size, aggregation conventions (arithmetic vs. geometric mean), and domain characteristics–—scientific abstracts are more formulaic than Wikipedia text. Therefore, a more meaningful metric is relative improvement.
Unsurprisingly, all LoRA-all models fine-tuned on datasets including hep-th abstracts (s1, s3-s8, s10) exhibited lower perplexity compared to the base model. On the contrary, models that did not include hep-th abstracts (s2, s9) had similar/higher perplexity. We also noted that the overall size of the dataset matters, with the lowest perplexity variants being those that included all hep-th samples augmented with other datasets (hep-ph, gr-qc, cs, q-bio).
@@
Llama $1.20^{+0.16}_{-0.13} 11$$1.20^{+0.16}_{-0.13}$
s1 $10.32^{+0.11}_{-0.09}$ $10.04^{+0.13}_{-0.10}$
s2 $11.38^{+0.11}_{-0.10}$ $10.62^{+0.15}_{-0.12}$
s3 $9.83^{+0.11}_{-0.09}$ $10.02^{+0.14}_{-0.11}$
s4 $10.41^{+0.11}_{-0.10}$ $10.08^{+0.12}_{-0.10}$
s5 $10.35^{+0.11}_{-0.09}$ $10.07^{+0.12}_{-0.10}$
s6 $10.34^{+0.09}_{-0.08}$ $10.05^{+0.13}_{-0.11}$
s7 $10.00^{+0.10}_{-0.09}$ $10.04^{+0.14}_{-0.12}$
s8 $9.90^{+0.08}_{-0.08}$ $10.04^{+0.12}_{-0.10}$
s9 $12.16^{+0.28}_{-0.19}$ $10.64^{+0.14}_{-0.11}$
s10 $9.93^{+0.11}_{-0.09}$ $10.00^{+0.15}_{-0.11}$
Table 3: Comparison of perplexity means across base and fine-tuned model configurations. Lower perplexity indicates better model performance. The perplexity of the base Llama model without any fine-tuning (LoRA-all or LoRA-QKV) has been included to make the comparison to the baseline easier. Uncertainties shown as $^{+\text{upper}}_{-\text{lower}}$ represent 95% confidence intervals from bootstrap resampling.
@@ 3.2 Semantic Similarity
On average we observed that the base model scored $0.88\pm 0.08$ while all the fine-tuned models (LoRA-QKV and LoRa-all) $0.90\pm 0.07$. Therefore, from the perspective of SemScore all models (base and fine-tuned) exhibit comparable performance in semantic similarity compared to the ground truth (with the fine-tuned models perhaps having a marginal edge). We note that the minimum values of the cosine similarities exhibited more pronounced variation: for the base model the minimum was 0.007, for the LoRA-all models the minimum was above 0.1 (with the exception of the s2 model, that was not trained on hep-th, at 0.06), and for the LoRA-QKV models the minimum had a much higher variation ranging from 0.015-0.015 to 0.15. For example, the s10 LoRA-QKV model (trained on hep-th, q-bio, cs) had minimum cosine similarity at 0.001-0.001.
@@
Category Mean SD Median NN
Base 3.75 1.52 4.0 36
LoRA-all 5.43 1.74 6.0 360
LoRA-QKV 4.92 1.70 5.0 360
Commercial 7.50 1.99 7.5 144
Table 4: Human evaluation scores by model category.
@@ 3.3 Human evaluation
To assess the quality of model-generated abstract completions, we conducted a human evaluation study with three domain experts, all experienced researchers in high-energy theoretical physics (hep-th). Each evaluator independently rated completions on a scale of 1–10 according to the following criteria:
  • 1–3 (Poor): Output is nonsensical, contains obvious errors, or is entirely generic with no connection to the specific physics context of the prompt.
  • 4–6 (Acceptable): Output demonstrates some domain awareness but may contain inaccuracies, lack specificity, or fail to coherently continue the abstract’s argument.
  • 7–9 (Good): Output is coherent, scientifically plausible, and appropriately continues the abstract with relevant physics content.
  • 10 (Excellent): Output is indistinguishable from a genuine hep-th abstract continuation, demonstrating deep domain understanding.
The evaluation corpus comprised 25 abstracts spanning diverse hep-th topics (string theory, conformal field theory, black holes, gauge/gravity duality), with completions generated by 25 model variants: the base Llama-3.1-8B model, 10 LoRA-all fine-tuned variants, 10 LoRA-QKV fine-tuned variants, and 4 commercial models (ChatGPT-4, Claude, Gemini, DeepSeek). This yielded 900 total human ratings, with 5 abstracts evaluated by all three raters, 1 by two and 19 individually. [^6] The human evaluation scores are also available as a CSV file in the GitHub repository https://github.com/Paul-Richmond/FeynTune.
Statistical Analysis:
Table 4 presents summary statistics by model category for all 900 ratings. Fine-tuned models significantly outperformed the base model (Mann-Whitney U=6,947U=6{,}947, $p<0.001$, rank-biserial $r=0.46$), while commercial models outperformed fine-tuned models (U=19,764U=19{,}764, $p<0.001$, $r=0.62$). The two LoRA variants showed a small but statistically significant difference (U=76,754U=76{,}754, $p<0.001$, $r=0.18$), with LoRA-all marginally outperforming LoRA-QKV.
Inter-Rater Reliability:
To assess evaluation consistency, we analyzed the 125 (abstract, model) pairs rated by all three evaluators. Pairwise Spearman correlations ranged from $\rho=0.47$ to $\rho=0.59$ (all $p<0.001$), indicating moderate agreement on relative rankings. The intraclass correlation coefficient ICC(2,1) =0.49=0.49 [95% CI: 0.30–0.64] reflects fair single-rater reliability, while ICC(2,k) =0.74=0.74 [95% CI: 0.56–0.84] indicates acceptable reliability when averaging across raters [24]. Agreement rates were 21% exact, 54% within $\pm 1$ point, and 77% within $\pm 2$ points—consistent with the inherent subjectivity of evaluating scientific text quality.
In addition to the above metrics, for which we give a qualitative description in Appendix E, we also report on some notable qualitative observations:
@@
The fine-tuned models could use technical language and, although training exclusively on abstracts limited their factual accuracy, demonstrated competitive performance relative to commercial LLMs. Table 5 demonstrates how two of our fine-tuned models (s3 and s8 LoRA-all) performed against the ground truth and ChatGPT-o. The fine-tuned models can clearly use technical language in sensible and relevant contexts, but are not 100% factually correct. For instance, in this example, the s3 model sensibly refers to the class-$\mathcal{S}$ $N=2$ superconformal field theories, which arise in compactifications of M5-branes, while the s8 model similarly picks up on the related AGT correspondence. The completion of ChatGPT, on the other hand, is very close to the content of the actual paper referencing key components of the proof. The completion of the base Llama model repeats elements of the prompt, as well as a line of (false) metadata at the end.
@@
In App. D we further discuss some of these observations with a collection of a few additional explicit completions.
@@ Appendix A Data Curation
$\frac{\text{size of hep-th}}{\text{size of hep-ph}+\text{size of gr-qc}}=0.5379232194539302$
@@
Appendix B Training Hyperparameters and Inference Settings
For full reproducibility, we provide complete specifications of all training hyperparameters and inference settings. These expand upon the summary provided in the main text, Table 2. Table 7 presents the complete training configuration used for fine-tuning the Llama-3.1-8B base model with LoRA. Table 8 presents the complete decoding configuration used during inference. Parameters not explicitly configured used HuggingFace Transformers library defaults (v4.45.2).
Parameter Value
Model Configuration
Base model meta-llama/Meta-Llama-3.1-8B
Attention mechanism Flash Attention 2
Quantization 4-bit NF4 with double quantization
Compute dtype bfloat16
LoRA Configuration
Rank (r) 8
Alpha 32
Dropout 0.05
Target modules q, k, v, o, up, down, gate projections
Training Configuration
Training epochs 4
Batch size (per device) 16
Gradient accumulation 1
Max gradient norm 1.0
Mixed precision bfloat16
Optimizer (AdamW)
Learning rate 3e-4
Betas [0.9, 0.95]
Epsilon 1e-5
Weight decay 0.1
Learning Rate Scheduler
Type Cosine with minimum LR
Warmup ratio 0.1
Minimum LR rate 0.01 (1% of initial LR)
Minimum LR (absolute) 3e-6
Dataset
Training data LLMsForHepth/hep-th_primary (train split)
Tokenizer meta-llama/Meta-Llama-3.1-8B
Random seed 42
Table 7: Training Hyperparameters
Parameter Value Source
Model Configuration
Attention mechanism SDPA Configured
Compute dtype float16 Configured
Generation Configuration
max_new_tokens 1024 Configured
min_new_tokens 1 Configured
temperature 0.7 Configured
do_sample true Configured
top_p 0.9 Default
top_k 50 Default
repetition_penalty 1.0 Default
no_repeat_ngram_size 0 Default
Stop criteria EOS token Default
Random seed 42 Configured
Padding side left Configured
Cache implementation static Configured
Table 8: Text Generation Parameters
Appendix C Cross Entropy Loss and Perplexity
@@
$\mathcal{L}=-\frac{1}{T}\sum_{t=0}^{T}\sum_{i\in|\mathcal{V}|}y_{t+1,i}\log(p_{\theta}(w_{t+1,i}|w_{1:t}))\;.$ (C.2)
@@
$\mathcal{P}=\Bigg(\prod_{t=0}^{T}\frac{1}{p_{\theta}(w_{t+1,\text{Id}(y_{t+1})}|w_{1:t})}\Bigg)^{1/T}\;,$ (C.3)
@@
The fact that all three ended up with the same perplexity of 3 is purely stochastic. Let us see what happens when we do a sentence completion with the context length set to two words:
@@
What we just discussed is the ground truth perplexity coming from the fact that the training data itself has ambiguity in how a sequence is completed having started off with "I". This is the lowest perplexity achievable. Models in general will have higher perplexity. The normalization is done so as able to compare perplexities across sequences of different lengths.
The above discussion involved the perplexity evaluated against one ground truth sequence. How do we compute it for a set of sequences? A naive answer, commonly used in the literature as it is computationally simpler, would be to take the cross entropy loss for the whole dataset and exponentiate it. This would amount to a geometric mean of perplexities of sequences. Another possibility is to take the arithmetic mean of perplexities of the sequences as this would give the average number of branching factors given a prompt.
@@
Appendix D Additional examples of completions
@@
Appendix E Evaluation Metrics
In this appendix we provide a qualitative description of the metrics used to assess the human evaluation part of this manuscript.
E.1 Statistical Comparison Metrics
  • Mann-Whitney U Test [26]: A non-parametric test that compares scores between two groups (e.g., fine-tuned vs. commercial models) to determine if one systematically scores higher than the other, without assuming that scores follow a Gaussian distribution. The test works by pooling all scores from both groups, ranking them, and computing the U statistic—essentially the number of times a score from group A exceeds a score from group B. For intuition: if group A consistently outperforms group B, most A scores will rank above most B scores, yielding a small $U$ (or equivalently, a large $U$ for the reverse comparison). The $U$ statistic alone is difficult to interpret directly; instead, one examines the $p$-value and effect size.
  • p-value: The probability of seeing a difference this large by chance alone. Values below 0.05 are conventionally considered statistically significant.
  • Effect Size (r): The rank-biserial correlation [27], measuring how large the difference is between groups on a scale from 0 to 1. This is a useful complement to the $p$-value, since large samples can yield small $p$-values even for negligible real-world differences. Interpretation: $|r|<0.3$ (small), $0.3\leq|r|<0.5$ (medium), $|r|\geq 0.5$ (large) [28].
E.2 Inter-Rater Reliability Metrics
  • Spearman $\rho$ [29]: Measures whether two raters rank completions similarly (e.g., both rank completion A above B), even if they use different parts of the 1–10 scale. Values near 1 indicate strong agreement on relative ordering.
  • ICC (Intraclass Correlation) [30]: Measures how consistently raters assign the same absolute scores, not just the same rankings. ICC(2,1) quantifies single-rater reliability; ICC(2,k) quantifies the reliability of the average across $k$ raters. Values above 0.75 are generally considered good, and above 0.5 moderate [31].
  • Agreement Rates: Measures the fraction of cases where raters gave the same score (exact), or scores differing by at most $\pm 1$ or $\pm 2$ points.